OpenStudy (anonymous):

2cos(3x-pi/4)=-sqrt3 find all solutions.

OpenStudy (anonymous):

\[2\cos(3x-\frac{ \pi }{ 4 })=-\sqrt{3}\]

OpenStudy (anonymous):

Start by dividing by 2.

OpenStudy (anonymous):

You get cos(3x-pi/4)=-sqrt3/(2)

OpenStudy (anonymous):

3x-pi/4=arccos(-sqrt3/2)

OpenStudy (anonymous):

You following?

OpenStudy (anonymous):

would I not use cos(A+B)=cosAcosB-sinAsinB?

OpenStudy (anonymous):

You could use that if you feel comfortable using that formula.